OpenStudy (anonymous):

What is the minimum value of the function: f(x)=2x^2-10x+5?

OpenStudy (mimi_x3):

Have you differentiated it?

OpenStudy (anonymous):

No I have not. How do you do that?

OpenStudy (mimi_x3):

do you know calculus?

OpenStudy (anonymous):

No I havent taken that class yet. This is AP Stats.

OpenStudy (mimi_x3):

Statistics..? This is Calculus. Or is there an algebra method to solve this?

OpenStudy (anonymous):

This class is AP Stats but U guess it is just basic math review. (Algebra)

OpenStudy (anonymous):

I mean I

OpenStudy (mimi_x3):

Hm, I don't know the algebra method.. maybe @lgbasallote might know

OpenStudy (lgbasallote):

i dont know algebra sorry

OpenStudy (unklerhaukus):

when the first derivative is zero, there will be a minimum or maximum

OpenStudy (mimi_x3):

is the algebra method got to do something with the vertex (-b)/2a?

OpenStudy (unklerhaukus):

yes Mimi_x3, that formula will give us the x-coordinate of the vertex,

OpenStudy (anonymous):

um i am not sure, but i think if I use (-b)/2a the answer would be 5 or 2.5?

OpenStudy (unklerhaukus):

the value is the f(x), so sub x=-b/2a into f f(-b/2a)= to get the value of f(x) that the minimum

OpenStudy (mimi_x3):

Vertex: \[\large \left(-\frac{b}{2a},a\left(-\frac{b}{2a}\right)^{2}+b\left(-\frac{b}{2a}\right)+C\right) \] a> 0 => its a min a<0 => its a max. however, the calculus method is more nicer :)

OpenStudy (anonymous):

I got 2.5.i hope its right. thank you for helping.